OpenStudy (bellezabrilla):

Help with x^2-11x+19=-5

OpenStudy (bellezabrilla):

@sshayer

OpenStudy (sshayer):

add 5 to both sides

OpenStudy (sshayer):

you solve simultaneously

OpenStudy (danettsep):

using that find the x

OpenStudy (bellezabrilla):

uh after adding the 5 to both sides I got x^2-11x+24

OpenStudy (danettsep):

there are two x

OpenStudy (danettsep):

find two factors of 24 that add up to 11

OpenStudy (sshayer):

either use quadratic formula or make factors

OpenStudy (danettsep):

-11 i mean

OpenStudy (danettsep):

i gave you the factoring, just find the x

OpenStudy (bellezabrilla):

yeah so I used the AM method and I got (x-8)(x-3)

OpenStudy (bellezabrilla):

wait so what about the x^2 ????

OpenStudy (sshayer):

-5=5=0 you get 0 on R.H.S.

OpenStudy (bellezabrilla):

ohhhh !!! I see.... so it would be (x-8)(x-3)=0 ?

OpenStudy (sshayer):

-11=-8-3 -8*-3=24 1*24=24

OpenStudy (sshayer):

correct.

OpenStudy (sshayer):

i am leaving now.

OpenStudy (bellezabrilla):

kk once again thxs for your help !! :)